OpenStudy (anonymous):

solve each system by elimination x-10y=13 -3x-5y=-4

OpenStudy (anonymous):

Stop posting all of your homework problems. At least explain in your question what specific part of the question you don't understand or know how to do.

OpenStudy (anonymous):

i culda swore this is for HELP not JUDGEMENT

OpenStudy (anonymous):

Other people want help too, and you are filling up the queue with your questions when other people want help. Plus I want to help you and you will not learn if you just have people give you the answer.

OpenStudy (anonymous):

Scale either equation such that the coefficient of one variable in one equation has the same absolute value as the same variable in the other equation. Add or subtract to "eliminate" that one variable. Solve for the remaining variable. Then return with the value you solved for and solve for the other variable.

OpenStudy (anonymous):

-2x-15y=5?

OpenStudy (anonymous):

Show your steps and we can help.

OpenStudy (anonymous):

Multiply the top equation by 3 to get 3x - 30y = 39 and then add that to the bottom equation to get the result -35y = 35

OpenStudy (anonymous):

why 3?

OpenStudy (anonymous):

Because the coefficient of x in the second equation is also 3. If you do that adding the equations eliminates the x leaving you with only y which you can solve.

OpenStudy (anonymous):

x-10y=13------------1 -3x-5y=-4-----------2 2*2, -6x-10y=-8----------3 1-3, x-10y-(-6x-10y)=13-(-8) x+6x=21 7x=21 x=3-----------------4 put 4 into 1, x-10y=13 3-10y=13 -10y=10 y=-1 therefore x=3, y=-1